Omicron Technologies has $50 million in excess cash and no debt. The firm expects to generate additional free cash flows of $40 million per year in subsequent years and will pay out these future free cash flows as regular dividends. Omicrons unlevered cost of capital is 10% and there are 10 million shares outstanding. Omicron's board is meeting to decide whether to pay out its $50 million in excess cash as a special dividend or to use it to repurchase shares of the firm's stock. Assume that you own 2500 shares of Omicron stock and that Omicron uses the entire $50 million to pay a special dividend. Suppose you are unhappy with Omicron's decision and would prefer that Omicron used the excess cash to repurchase shares (you will not sell in case of a repurchase). The number of shares that you would have to buy in order to undo the special cash dividend that Omicron paid is closest to:

125

275

250

310
Answer #1

Enterprise value = $40 million / 10% = $400

Market value = 400+ 50 = 450 miliion

Market value per share = 450 million / 10 million shares = $45 per share

Special dividend per share = $50 million / 10million shares = $ 5 per share

Dividend per share = 2500 shares * $5 = 12500

No of shares to be bought = 12500 / 45 = 277.78 closest to 275 shares.

Option 275 shares.